# AI Rules? Characterizing Reddit Community Policies Towards AI-Generated Content

### v1

@@ 4.3. RQ3: What Types of Rules Exist for Governing AI?
Text 5.23 This includes both AI generated images/artwork and text generated by tools such as ChatGPT.
@@
Deepfake 5.08 Deepfakes or other types of fakes are not allowed here, you will be banned for posting them.
@@
Bots 3.37 Karma-farming with ChatGPT bots, Malware, phishing, spam or scam will result in your account being banned without warning.
Videos 2.84 No low effort AI content or Videos.
Framing Generate 51.49 AI-generated content of any kind is forbidden.
@@
Assist 23.69 Traditional paintings and drawings only. No videos, screenshots, photography, audio, multi-panel comics, content generated or AI-assisted submissions.
Unspecified 8.07 The removal of low quality and, AI submissions, are subject to remove at mod discretion.
@@
Since the Art and Celebrity Topics featured prominently in Section 4.2, we also separately examined patterns among Rule Types within each of these Topics. Among subreddits in the Art Topic, we find a higher share of AI rules with the Copyright and Piracy label ($34\%$). Among Celebrity subreddits, we see that the Low Quality Content and Format Rule Types labels are much more common ($74\%$ and $71\%$). In terms of the explicit rationales given for the AI rules, almost half of all AI rules in Celebrity subreddits were labeled Rationale: Low Quality and $29\%$ were labeled Rationle: Inauthentic. Additionally, $86\%$ of AI rules from Celebrity subreddits explicitly govern the use of “Fakes” (labeled with Medium: Fakes), which upon closer examination appears to be referring to heavily-edited photos.
@@ 5. Discussion
Our quantitative study offers two main research contributions: (1) a large-scale analysis of rules about AI use in Reddit communities and (2) a taxonomy for classifying these rules. Our study is an effort to support community self-moderation inspired by Seering’s second “guiding question” for the HCI research community: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” (Seering, 2020). The sudden arrival of generative AI poses unique challenges for online communities (Lloyd et al., 2023) that HCI research is uniquely positioned to help address. Below we put our findings in dialogue with prior work and discuss design implications for social platforms.
Our findings show that subreddit communities, especially large ones (in terms of number of subscribers), are changing their rules to respond to generative AI. The new rules focus on AI-produced images more than any other medium, and often cite concerns about quality. The likelihood of having rules about AI varies with the type of community, with communities related to art or celebrities more likely to have these rules, and communities that are oriented around social support less likely.
@@
The differences between our results and those of Fiesler et al. (Fiesler et al., 2018) emphasize the importance of governing AI usage to image-sharing communities. Our findings that Restrictive AI rules are more common than Prescriptive ones seems to suggest that, at least at this point in time, communities appear more interested in clarifying how AI cannot be used than how it can. Still, there is some evidence of flexibility in these stances, and it will be interesting to observe if they grow more or less flexible over time. We reiterate Fiesler et al.’s point that rules on Reddit are context-sensitive, so it is hard to generalize about the situations when communities opt for disclosure rather than an outright ban, but digging into these cases could be a topic of future HCI research.
Evaluating our results in the context of the qualitative findings of Lloyd et al. (2023) paints a nuanced picture of the prevalence of some of the concerns raised in their interviews. For example, over $50\%$ of rules in our AI Rules Set are about images, but the majority of concerns raised in Lloyd et al. pertain primarily to text. This is a significant difference as helping communities respond to AI-generated text as opposed to AI-generated images could be a very different task. Another interesting difference is that we find the Social Support Community Archetype (Prinster et al., 2024) to be negatively correlated with the presence of AI rules, which goes against one of the theories presented by Lloyd et al., that communities who seek to encourage authentic human connection may object the most to AIGC. We do find concerns about authenticity, but it tends to be more about the authenticity of content, not of interpersonal communication. Our findings that Copyright and Piracy AI rules are more common in Art subreddits echoes some of the quotes from the moderators’ of art-related communities interviewed in Lloyd et al. who object to AI because of how current AI models have been trained. The prevalence of AI rules in Celebrity subreddits also echoes the concerns about authenticity raised in Lloyd et al.’s interviews. Rules from Celebrity subreddits tended to be about deepfakes or heavily altered images, which recent studies of AIGC in the wild have identified as the most common form of misuse (Marchal et al., 2024). Lloyd et al. also emphasizes a concern that AIGC is low-quality content, which is confirmed by our analysis to be the most universal explicit objection across the platform. What a community considers to be “low quality” is extremely context-dependent, but this could be one area where UI features that clearly communicate norms and expectations to community members, such as rules that display at post creation time, could go a long way to positively shape user behavior (Matias, 2019).
Looking ahead, it will be interesting to see how these rules—and norms—continue to change over time. So far our data shows that AI rules have become more prevalent, but it is possible that these rules could disappear if AI use becomes widespread to the point that it is unremarkable, or where bans are impossible to enforce. Communities who care about gate-keeping (Keegan and Gergle, 2010) may seek out other ways for users to demonstrate their authenticity, similar to those that have developed in some anonymous online communities (Bernstein et al., 2011). The HCI community should keep an eye on how these rules change over time in order to continue to support these communities as their norms and practices evolve.
Our study is limited in that it is based on a sample of available metadata for subreddits that were publicly listed on Reddit’s web interface in the summer of 2023. We recollected metadata for this sample one year later, but did not add any new subreddits, which means that new communities created since our initial collection are excluded from our study. By collecting only publicly listed subreddits for our sample, our sample is also missing subreddits that are public but unlisted. This makes it likely that certain types of subreddit are underrepresented, such as NSFW subreddits. Our sample also excluded subreddits with a primary language other than English. Future research could explore whether the trends that we observed still hold across the types of communities excluded from our sample.
@@
Our study also limits our analysis to text that is contained within a subreddit’s ”rules” metadata field. There are various other ways that online communities communicate rules and norms, such as through public discussions, wikis, direct messages, or moderation actions. Future work interested in studying communities’ attitudes towards AI use could consider these other sources of data, as well as qualitative approaches that would elicit direct input from community members and moderators.
@@
Ethics Statement. Our study was not required to undergo IRB review as we worked only with publicly available data. Still, we recognize the importance of conducting ethical social media research and have taken steps to follow the best practices of the HCI community (Fiesler et al., 2024). We did not collect any information about individual users or conversations. Additionally, we did not include in our study any subreddits who were not listed in Reddit’s public index of communities. We removed from our sample any subreddits who were public during our first crawl but no longer accessible during our second. Our paper does not reference any of the specific communities in our sample, only aggregate statistics. We plan to make our datasets public after this publication completes the peer-review process, but will take additional steps to ensure no sensitive information is contained before doing so.
@@ 6. Conclusion
Reddit communities across the platform have responded to the arrival of generative AI by updating their public rules to state explicit positions on the use of this new technology. A specific community’s attitude towards the use of AIGC is nuanced and context-dependent, but our findings suggest a general trend of opposition from communities who value quality and authenticity. It remains to be seen how these rules will be enforced, but HCI researchers should stay attuned to the changing needs and norms of these online communities as they continue to adapt to this new technology.
@@ A.1. Detecting Community Rules Governing AI Use
A.2. Applying Subreddit Classification Labels
2You will be given a JSON string containing metadata about a subreddit that will contains its name, topic, public description, description and rules.
5
6Category: Community Archetype
8Labels:
9- Name: Topical Question and Answer
11- Name: Learning and Perspective Broadening
13- Name: Social Support
15- Name: Content Generation
17- Affiliation with an Entity
19
20Category: Topic
22Labels:
23- Video Games
24- Image Sharing
25- Entertainment
26- Personal
27- Technology and Science
28- Celebrity
29- Hobby
30- Local
31- Sports
32- Support
33- Music
34- Art
35- Humor
36- Meta
37- Business and Organizations
38- Animals
39- Work
40- Writing
41- Learning
42- Politics
43- Games
44- News
45- Health
46- Food
47- Parody
48- Culture
49- Fashion
50- Drugs
A.3. Applying Rule Classification Labels
@@
79Category: Framing
80Label name: framing--generate
@@
83Label name: framing--assist
@@
86Label name: framing--unspecified
@@
124Label name: medium--text
@@
127Label name: medium--images
@@
130Label name: medium--videos
@@
133Label name: medium--art
@@
136Label name: medium--deepfake
@@
139Label name: medium--fakes
@@
142Label name: medium--bots
@@
145Label name: medium--content
@@
148Label name: medium--unspecified
@@
153Instructions: Does this rule ban or disallow AI or some sort of content made using AI without providing exceptions?

### v2

@@ 4.3. RQ3: What Types of Rules Exist for Governing AI?
Text 6.59 This includes both AI generated images/artwork and text generated by tools such as ChatGPT.
@@
Deepfake 4.6 Deepfakes or other types of fakes are not allowed here, you will be banned for posting them.
@@
Bots 3.34 Karma-farming with ChatGPT bots, Malware, phishing, spam or scam will result in your account being banned without warning.
Videos 3.95 No low effort AI content or Videos.
Framing Generate 53.88 AI-generated content of any kind is forbidden.
@@
Assist 16.08 Traditional paintings and drawings only. No videos, screenshots, photography, audio, multi-panel comics, content generated or AI-assisted submissions.
Unspecified 12.02 The removal of low quality and, AI submissions, are subject to remove at mod discretion.
@@
Since the Art and Celebrity Topics featured prominently in Section 4.2, we also separately examined patterns among Rule Types within each of these Topics. Among subreddits in the Art Topic, we find a higher share of AI rules with the Copyright and Piracy label ($25.54\%$). These Copyright and Piracy rules primarily focus on proper attribution and posting original content. Among Celebrity subreddits, we see that the Low Quality Content and Format Rule Type labels are much more common ($58.32\%$ and $55.32\%$). The Format label means that references to AI often co-occur with regulations about the format of posts — such as statements about how posts should be titled or specifications about minimum word counts. In terms of the explicit rationales given for the AI rules, almost half of all AI rules in Celebrity subreddits were labeled Rationale: Inauthentic and $34.79\%$ were labeled Rationale: Low Quality. Additionally, $77\%$ of AI rules from Celebrity subreddits explicitly govern the use of “Fakes” (labeled with Medium: Fakes), which upon closer examination appears to be referring to heavily-edited photos (these rules separately make references to “Deepfakes”, which we distinguish with a different label).
@@ 5. Discussion
Our quantitative study offers three main research contributions: (1) a large-scale analysis of rules about AI use in Reddit communities, (2) a taxonomy for classifying these rules, and (3) a dataset of subreddits, their metadata, and public community rules at two distinct points in time. Our study is an effort to support community self-moderation inspired by Seering’s second “guiding question” for the HCI research community: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” (Seering, 2020). The sudden arrival of generative AI poses unique challenges for online communities (Lloyd et al., 2023) that HCI research is uniquely positioned to help address. Our study attempts to measure at scale the prevalence and variety of rules that online communities have enacted in response to this new technology. We hope that our data and analysis can add additional context to qualitative studies into online communities’ responses to AIGC, as well as to general knowledge about online community self-moderation practices. Below we put our findings in dialogue with prior work and discuss design implications for social platforms.
5.1. The Role of Community Rules
While we focus our analysis on rules about AI use, one of our primary contributions is a dataset of community rules that the HCI research community can use to answer a wider range of questions about community self-moderation. As far as we know, this is the first dataset of its kind, as prior studies of self-governing online communities’ rules (Fiesler et al., 2018; Nicholson et al., 2023; Cai et al., 2021; Hwang and Shaw, 2022; Reddy and Chandrasekharan, 2023) have not made their rules datasets public. Large-scale, quantitative analysis of these rules offers a view into platform-wide patterns that complement past qualitative HCI research into online community self-moderation. Our AIGC-focused analysis finds that subreddit communities, especially those with more subscribers, are changing their rules to respond to generative AI. But what do these rules tell us about the experience of participating in or moderating these communities?
Past research into community self-moderation has found that community rule changes are most often enacted, “in response to unexpected incidents” that impact a community (Seering et al., 2019). Qualitative work on the experience of moderating AIGC (Lloyd et al., 2023) provides evidence that AI rules are created for the same reason, as many moderators report problematic AI use in their communities as a precursor to their formal rules. This mechanism of rule creation suggests that a community’s rules are not a perfect proxy for its values, as community values do not need to be codified into rules unless there is some incident that surfaces community misalignment. Seen in this light, the communities with AI rules in our dataset are not necessarily the communities with the strongest stances on AI, but those where some misalignment about AI use emerged between community members. Viewing rules as evidence of past misalignment could explain why Social Support subreddits were negatively associated with AI rules in our analysis: perhaps these communities are more aligned in general, so that even if they have strong stances on AIGC, they do not need rules to enforce them.
It is worth noting that while rules may indicate some past community misalignment, they do not necessarily tell us how rules shape behavior in practice. For example, prior work has shown that humans cannot reliably detect AIGC (Jakesch et al., 2023), which raises the question of how enforceable AIGC rules are. At the same time, it is unclear whether or not AIGC rules even need to be enforced: past work on community rules suggests that they can shape behavior by clarifying norms (Matias, 2019; Lloyd et al., 2023), even without an effective enforcement mechanism. Still, it is reasonable to assume that some communities with AI rules are attempting to enforce them. We see the prevalence of AI rules, especially in the largest subreddits, as evidence that HCI research should explore design solutions that can help with their enforcement.
One way that HCI research can empower communities to effectively enforce their AIGC policies is by offering tools for policy design. While some online communities are enacting top-down AI policies (Overflow, [n. d.]), Reddit has so far left this decision to individual communities. We like this approach, as past research has shown that community moderators learn by doing and interacting closely with their community (Cullen and Kairam, 2022), resulting in a quick feedback loop that better positions them to craft context-sensitive policies (Seering, 2020). Still, opportunities exist for HCI research to help communities enact effective policies, both by offering empirical evidence of the effects of certain types of rules, as well as new systems for policy development (Zhang et al., 2020). Future work could use our rules dataset to find similar communities that differ in their policy choices, which could be used as a natural experiment to study the effects of different rules. Studying the effects of policies may uncover general knowledge about which kinds of rules (such as “prescriptive” vs “restrictive”) encourage community engagement and which cause members to leave for other communities. Platforms could use this knowledge to offer moderators a reasonable set of “default” rules at community creation time, based on the type of community being created. Some alternative systems for policy development from past HCI research have explored ways of engaging more members of a community, beyond just moderators, in the policy creation process (Zhang et al., 2020). We see these more democratic systems of rule development as promising at a time when norms around AI use are still in flux.
5.2. Supporting Community Self-Governance in the Age of Generative AI
@@
Comparing our findings to those from an earlier large-scale analysis of Reddit rules conducted by Fiesler et al. (Fiesler et al., 2018) provides insight into what is different about AI rules. We found that rules governing AI usage are more likely to be about images than the general rules analyzed in Fiesler et al. Together with the finding that images are the medium most referenced in AI rules, this finding suggests that the use of AI images may be the most common cause of community misalignment about AI norms. Our findings that Restrictive AI rules are more common than Prescriptive ones seems to suggest that, at least at this point in time, communities appear more interested in clarifying how AI cannot be used than how it can. Still, there is some evidence of flexibility in these stances, and it will be interesting to observe if community stances towards AI grow more or less flexible over time. It may be difficult to generalize about when communities require AI disclosure vs when they ban it outright, but digging into the motivations for these different approaches could be a topic of future HCI research.
These rules—and norms—will continue to change over time. So far our data shows that AI rules have become more prevalent, but it is possible that these rules disappear if AI use becomes widespread to the point that it is unremarkable, or where bans have negative impacts on community engagement. Two of the top three most common rationales in the AI rules in our data were “Inauthentic” and “Low effort”, which seems to suggest that some communities perceive that AI use cheapens the social interactions that occur in online communities. These rationales for AI rules echo the findings from prior studies into the use of automated tools in interpersonal communication (Moradi and Levy, 2024; Rae, 2024). We suspect that even if communities cannot effectively ban AIGC, those that care about gate-keeping (Keegan and Gergle, 2010) may seek out other ways for users to demonstrate their authenticity, similar to those that have developed in some anonymous online communities (Bernstein et al., 2011). Platform designers can help by offering communication tools that encourage authenticity, such as those explored in recent work on effortful communication (Kelly et al., 2017; Zhang et al., 2022). Additionally, past HCI studies of online communities have shown that close-knit communities with strong social identity are able to navigate the challenging task of developing new norms around controversial technology use (Fiesler and Bruckman, 2019). In the age of generative AI, HCI research could develop novel ways to encourage strong social identity formation within online communities.
Our results complement Lloyd et al. (2023)’s qualitative investigation into moderators’ experience of enforcing AIGC rules by documenting the prevalence of these rules across Reddit. Their study identifies three main categories of moderator concern about how AIGC will impact their communities: concerns about content quality, social dynamics, and governance processes. Of these, content quality concerns are the most represented in our data. What a community considers to be “low quality” is extremely context-dependent, but this could be one area where UI features that clearly communicate norms and expectations to community members, such as rules that display at post creation time, could go a long way to positively shape user behavior (Matias, 2019). We suspect that rules about social dynamics are less prevalent in our data because of the phenomenon that we discuss in Section 5.1, where desired social dynamics may not need to be codified into rules because communities that care about them may be naturally better aligned. For example, in our data we do find rules about authenticity, but it tends to be more about the authenticity of content, not of interpersonal communication. Additionally, our findings provide further evidence of Lloyd et al.’s finding that moderators’ concerns about AIGC are based on their communities’ values. While rules usually do not explicitly state a community’s values, sometimes we can infer them, e.g. it is not a coincidence that Copyright and Piracy AI rules are more common in Art subreddits. Finally, we note that rules about images are very common in our data, while the majority of concerns raised in Lloyd et al. pertain primarily to text. This is a significant difference for the HCI research community to note, as helping communities respond to AI-generated text as opposed to AI-generated images is a different challenge with different potential solutions. For example, images may be more likely to be robustly watermarked (Dathathri et al., 2024) or to have provenance metadata (Feng et al., 2023), which platforms could clearly communicate to viewers.
@@
Our study is limited in that it is based on samples of subreddits that were publicly listed on Reddit’s web interface in November 2024 and July 2023. Our sample is thus missing subreddits that are public but unlisted, which makes it likely that certain types of subreddit are underrepresented, such as NSFW subreddits. Our sample also excluded subreddits with a primary language other than English. Future research could explore whether the trends that we observed still hold across the types of communities excluded from our sample. Our study also limits our analysis to text that is contained within a subreddit’s ”rules” metadata field. There are various other ways that online communities communicate rules and norms, such as through public discussions, wikis, direct messages, or moderation actions. Future work interested in studying communities’ attitudes towards AI use could consider these other sources of data, as well as qualitative approaches that would elicit direct input from community members and moderators.
@@
Ethics Statement. Our study was not required to undergo IRB review as we worked only with publicly available data. Still, we recognize the importance of conducting ethical social media research and have taken steps to follow the best practices of the HCI community (Fiesler et al., 2024). We did not collect any information about individual users or conversations. Additionally, we did not include in our study any subreddits who were not listed in Reddit’s public index of communities. We plan to make our datasets public after this publication completes the peer-review process, but will take additional steps to ensure no sensitive information is contained before doing so.
@@ 6. Conclusion
Communities across Reddit have responded to the arrival of generative AI by updating their public community rules with explicit positions on the use of this new technology. A specific community’s attitude towards the use of AIGC is nuanced and context-dependent, but our findings suggest a general trend of opposition from content generation communities that value quality and authenticity. It remains to be seen how these rules will be enforced, but HCI researchers should stay attuned to the changing needs and norms of these online communities as they continue to adapt to this new technology. Especially as social platforms begin incorporating generative AI features, it is important for designers to consider how to encourage the self-determination of communities with nuanced and varied stances on the use of AIGC.
@@ A.1. Detecting Community Rules Governing AI Use
A.2. Applying Rule Classification Labels
@@
79Category: Usage
80Label name: usage--generate
@@
83Label name: usage--assist
@@
86Label name: usage--general
@@
124Label name: form--text
@@
127Label name: form--images
@@
130Label name: form--videos
@@
133Label name: form--art
@@
136Label name: form--deepfake
@@
139Label name: form--fakes
@@
142Label name: form--bots
@@
145Label name: form--content
@@
148Label name: form--general
@@
153Instructions: Does this rule ban or disallow AI or some sort og content made using AI without providing exceptions?
@@
A.3. Applying Subreddit Classification Labels
2You will be given a JSON string containing metadata about a subreddit that will contains its name, public description, description and rules.
5
6Category: Community Archetype
8Labels:
9- Name: Topical Question and Answer
11- Name: Learning and Perspective Broadening
13- Name: Social Support
15- Name: Content Generation
17- Affiliation with an Entity
19
20Category: Topic
22Labels:
23- Video Games
24- Image Sharing
25- Entertainment
26- Personal
27- Technology and Science
28- Celebrity
29- Hobby
30- Local
31- Sports
32- Support
33- Music
34- Art
35- Humor
36- Meta
37- Business and Organizations
38- Animals
39- Work
40- Writing
41- Learning
42- Politics
43- Games
44- News
45- Health
46- Food
47- Parody
48- Culture
49- Fashion
50- Drugs